OpenStudy (anonymous):

fine the domain of f(x)^-1 if f(x) = 4x-8

OpenStudy (accessdenied):

i want to assume this is the inverse of f, but juts to make sure... is this the inverse of f or 1/f(x)?

OpenStudy (anonymous):

f(x)

OpenStudy (anonymous):

Looks like they want an inverse function.

OpenStudy (anonymous):

y = 4x-8 Solve for x. (y-8)/4 = x y/4 - 2 = x Change the y to x, and change the x to y. y = x/4 -2 f(x)-1 = x/4-2 Can someone verify this please?

OpenStudy (anonymous):

I've gotten much better at these!

OpenStudy (anonymous):

These are answer choices x≠ 0 x≠ 1/4 x≠ -4 All real numbers

OpenStudy (anonymous):

Well my guess would be all real numbers for the inverse function. What do the rest of you guys think?

OpenStudy (anonymous):

both the function and the inverse are lines. so yes, all real numbers...

OpenStudy (anonymous):

@Lib Libs "I've gotten much better at these!" Awesome, I'm glad to hear it! Ex: 1 inverse functions: http://tutorial.math.lamar.edu/Classes/Alg/InverseFunctions.aspx

OpenStudy (anonymous):

I'm working on these and the one's with fractions :/

OpenStudy (anonymous):

got another one you're stumped on?

OpenStudy (anonymous):

Let f(x) = x^2 + 6 and g(x) = x + 8 over x find (g of f)(4)

OpenStudy (anonymous):

g(x^2+6) = (x^2+6) +8 =x^2 +14 g(f(4)) = (4)^2 +14 = 30

OpenStudy (anonymous):

55/63 15/11 11/15 63/55 Are answer choices

OpenStudy (anonymous):

Does this problem have to do with inverses? Do you have a screenshot?

OpenStudy (anonymous):

Yes inverse's http://i1204.photobucket.com/albums/bb402/Libby_Williams/Yoyoyo.jpg

OpenStudy (anonymous):

Weird b/c it doesn't mention anything about inverses. But I'll just solve it that way and see if I stumble on something.

OpenStudy (anonymous):

Maybe its not

OpenStudy (accessdenied):

nope, its just function composition

OpenStudy (anonymous):

Yes thats what it is?

OpenStudy (anonymous):

@AccessDenied did you come up with a different answer from what I got?

OpenStudy (anonymous):

Sorry ; - ;

OpenStudy (accessdenied):

i dont think you took into account the (x+8) / x, like i did the first time i went through it and got the same answer

OpenStudy (anonymous):

I didn't see the "over x." I wasn't paying attention lol.

OpenStudy (accessdenied):

i didnt catch it either, and i was confused why it wasnt in the answers. :P it does seem like it goes to be 22+8 / 22, which is 30/22 or 15/11

OpenStudy (anonymous):

Let f(x) = x^2 + 6 and g(x) = x + 8 over x find (g of f)(4). g(f(4)): I'll just find f(4) and then plug it into g. f(4) = 16 +6 = 22 g(22) = (22+8)/22 = 30/22 =15/11